OpenStudy (anonymous):

Could someone give me some tips for this kind of problem? Determine a region of the xy-plane such that the given DE has an unique solution whose graph go through the point (x0, y0) in this region.

OpenStudy (anonymous):

\[ \frac{dy}{dx} = y^{2/3} \]or \[ \frac{dy}{dx} = \sqrt{xy} \]

OpenStudy (anonymous):

i'm sure it's DG-Universe lol

OpenStudy (anonymous):

I know this shouldn't be exactly hard, but I don't know how to do it :-(

OpenStudy (turingtest):

*bookmark

OpenStudy (anonymous):

I don't have the answer for the second one. For the first one, it should be: semi-plan defined by y > 0 or y < 0.

OpenStudy (anonymous):

Maybe solving the DE and checking where the solution is defined? In the existence and uniqueness theorem, we consider a rectangular region. So, I will try that.

OpenStudy (amistre64):

it looks like xo,yo satisfies an initial condition right?

OpenStudy (anonymous):

I thought so too, but all we are given about x0 y0 is that. I guess we have to solve for x0?

OpenStudy (amistre64):

are these diffyQs on the same xy plane? or two different questions altogether?

OpenStudy (anonymous):

two different questions.

OpenStudy (anonymous):

I want help in any of these, so I can understand how to solve these kind of problem.

OpenStudy (anonymous):

this* kind

OpenStudy (amistre64):

then id assume that the solution is an interval on which the initial conditions are satisfied theres no direction to make it a largest interval so any suitable one should suffice

OpenStudy (amistre64):

the eqs seem rather separable to get a solution from

OpenStudy (anonymous):

Hmm. You mean, for the first one my solution would be: x + c = 3y^(1/3). Then, I should pick > 0 or < 0 for this to hold?

OpenStudy (anonymous):

I mean, if we pick x = 0, y = c, but I don't get why we can only pick y > 0 or y < 0

OpenStudy (amistre64):

uniqueness implies that in the interval there is only one option that can result. right? without knowing more about the problem, that is my best idea

OpenStudy (anonymous):

That's still confusing for me. And Boyce did not make this any simpler. Should I always disconsider the zero of the solution? Say, \( (4 - y^2)y ' = x^2 \). We have to pick y > 2 or y < -2 or -2 < y < 2.

OpenStudy (anonymous):

Because if it's zero, we have infinite solutions?

OpenStudy (amistre64):

the xo and yo conditions help to narrow down an interval, but so far i havent been able to interpret your conclusions. 4y - y^3/3 = x^3/3 + C , using xo and yo to calculate for C we can define a suitable y for this. as such, I see no restrictions that would chop up the function into different intervals: no vertical asymptotes, no holes, no undefineds ....

OpenStudy (anonymous):

Still, the answer provided by the book is the one I said above. :-( That's why it's so confusing for me.

OpenStudy (amistre64):

maybe the explicit form has some issues tho

OpenStudy (amistre64):

notice that their y intervals chop up the graph into "functions" on the intervals

OpenStudy (amistre64):

a solution has to be a function as well

OpenStudy (anonymous):

So, I have to "graph" the solution and see where it's a function and if it's continuous (and if its y derivative is also continuous?)